DeezNutzLmao:

Carlos and his family are driving to California for vacation. The trip is 1,146 miles and they drive at an average speed of 56 mph. Which equation would give the number of miles remaining until they reach their destination, M, in terms of h, the number of hours they have driven?

DeezNutzLmao:

A. M = 56 - 1,146h B. M = 1,146 + 56h C. M = 56 + 1,146h D. M = 1,146 - 56h

ac307:

nice user

DeezNutzLmao:

@ac307 wrote:
nice user
thx lol

ac307:

also your answer is D

DeezNutzLmao:

tyty

jhonyy9:

@ac307 wrote:
also your answer is D
please explain your answer ,how you get this result @smokeybrown

SmokeyBrown:

@ac307 wrote:
also your answer is D
I agree with your answer, but as jhonny9 said, you should explain your reasoning instead of only giving the direct answer In this case, the total distance is 1146 at the beginning, and each hour the family travels 56 miles, so the remaining distance decreases by 56 miles. To calculate how many miles are left, we could subtract (56*h) from the total of 1146, which is what is written in equation D.